OpenStudy (anonymous):

Find an equation of the line containing the given pair of points. Express your answer in the form x=a, y=b, or y=mx+b (0,0) and (7,4) y=?

OpenStudy (anonymous):

find slope m=4-0 / (7-0) and use the given y=mx+b formula

OpenStudy (anonymous):

or y-y1=m(x-x1)